OpenStudy (anonymous):

lim x->-inf 1/2^x is the answer inf?

OpenStudy (bahrom7893):

no it's still 0

OpenStudy (anonymous):

how come lim x-> -inf e^-x is inf

OpenStudy (bahrom7893):

because in your case u're blowing up the denominator of a fraction... and the bigger it gets, the smaller the fraction gets.. but e^(-x) as x goes to -inf is e^(-(-inf))=e^inf=inf

OpenStudy (anonymous):

I have just put this limit(1/2^x, x = -infinity) in maple and the answer has come infinity

OpenStudy (anonymous):

http://www.numberempire.com/limitcalculator.php